# ProteinWeaver: A Divide-and-Assembly Approach for Protein Backbone Design

### v1

@@ 2.1 Divide and assembly diffusion framework
This spliced distance map $\bar{\mathbf{M}}$ initializes the edge representation, providing condition for the diffusion model, as shown in Appendix A.2.1. To enhance the spatial flexibility of domain assembly, we insert a 15-residue linker between adjacent domains and set its interaction with other resions to $\mathbf{-1}$. The model is tasked with learning the deformations generated by each domain in $\{\bar{\mathbf{S}}_{{D}_{1}},\bar{\mathbf{S}}_{{D}_{2}},...,\bar{\mathbf{S}}_{{D}_{m}}\}$ during the combination process as well as the patterns of the ultimate interactions.
To assemble generated domains, we utilize FrameDiff  (Yim et al., 2023), the $\mathrm{SE(3)}$ diffusion model to pretrain the domain assembly model. As illustrated in the Appendix A.2.1, the assembly module contains a series of folding blocks, where each folding block processes node representation, edge representation and structural frames.
@@ 2.2 Training
In this section, we provide the details of ProteinWeaver training, including datasets, pretraining, and preference alignments. More details can be found in Appendix A.2.1.
@@
Here, $\mathcal{L}_{\text{trans}}$ computes the loss between predicted translations with reference translations, while $\mathcal{L}_{\text{rot}}$ calculates the loss of rotation scores. $\mathcal{L}_{\text{atom}}$ represents the atom coordinate loss and $\mathcal{L}_{\text{pairwise}}$ computes the pairwise distance loss between predicted and reference positions. Following FrameDiff (Yim et al., 2023), we apply auxiliary losses when $t\leq{0.25}$. More details can be found in Appendix A.2.2.
@@ 3.1 Experimental setups
We evaluate ProteinWeaver’s performance using two tasks: (1) Domain assembly: this is a new protein backbone design task, introduced in our study, lacking existing deep learning-based baselines. It serves both as an appropriate scenario to evaluate our approach and as an important protein engineering task previously studied using traditional method (Huddy et al., 2024). (2) Backbone design: This task focuses on the design of de novo proteins, a subject of significant interest in the global research community. When generating backbone, we use the best of 3 operation (bo3). More details are provided in Appendix A.6.
@@
We pretrained the diffusion model for domain assembly using 5,835 filtered multi-domain PDBs from the RCSB database (Berman et al., 2000). For the preference alignment, we performed pairwise combinations of 100 single-domain structures, generating 10,000 pairs to build winner and loser datasets for alignment. Detailed methodology is provided in Appendix A.1.1 and A.1.2.
@@
We evaluate generation conformations as to their overall backbone quality, interface quality, novelty, and diversity. We also provide interface scTM and scRMSD metrics for domain assembly quality evaluation. More details are provided in Appendix A.5.
@@ 4.1 Diffusion modeling on protein backbone
Inspired by the considerable success of diffusion models across various fields (Ho et al., 2020; Song et al., 2020), researchers have applied this approach to protein structure and sequence design. Anand et al. pioneered this effort with a co-diffusion model for backbone, sequence, and sidechain generation using AlphaFold2’s Structure Module  (Anand & Achim, 2022) . Subsequent methods explored diffusion on inter-residue geometry  (Lee et al., 2023) and backbone dihedral angles  (Wu et al., 2024a). Current protein structure diffusion models primarily focus on end-to-end structure generation in SE3 or R3 space  (Yim et al., 2023), with extensions to function motif scaffolding  (Trippe et al., 2022; Yim et al., 2024). Chroma achieved more general conditioning and improved efficiency through geometrically constrained harmonic constraints  (Ingraham et al., 2023), while RFdiffusion demonstrated state-of-the-art designability with experimental validation  (Watson et al., 2023). However, despite these advancements, the capacity of these state-of-the-art methods for long-chain backbone generation remains limited, highlighting an area for further research and development in the field of protein design.
@@ Conclusions.
Despite these advances, several limitations of our current model warrant further investigation. (1) Controllability: The current version of ProteinWeaver lacks control over specific structural features or functional properties of the generated proteins. (2) Side-chain Information: Our model focuses on backbone design and does not include detailed side-chain information, which is crucial for precise protein function prediction and design. (3) Robustness: While ProteinWeaver shows promising results, the robustness of the domain assembly, especially the randomly sampled CATH domain, requires further improvement. Addressing these limitations in future work could significantly enhance the applicability and reliability of ProteinWeaver for diverse protein design tasks. (4) Spatial flexibility in domain assembly: In this study, we design protein backbones by assembling domains in one single chain. While this goal is effective in our current study, it may limit the exploration of more diverse spatial arrangements. Future research could investigate variable linker lengths or even linker-free assembly methods, to potentially address more challenging problems, such as protein-protein docking.
While our framework demonstrates a new paradigm for protein backbone design, we acknowledge several technical aspects that warrant further investigation: (1) Recent advancements have revealed that flow matching-based protein backbone generation methods outperform diffusion-based approaches. An exciting avenue for future research lies in exploring whether replacing our current diffusion-based framework with a flow matching-based approach could yield further benefits; (2) Domain structure representation: Our current method explicitly represents local domain structures as distance maps. However, the efficacy of alternative pairwise representation methods remains an open question. Investigating various representation techniques could provide crucial insights into the most effective ways to capture and utilize structural information in protein design, potentially uncovering superior approaches. (3) Integration of design stages: Currently, our approach employs two separate steps for backbone design. A promising direction for future research is developing an end-to-end unified framework that efficiently integrates these steps. Future research could further refine and extend this framework by addressing these questions, potentially revolutionizing our ability to create functional protein backbones with unprecedented precision.
Appendix A Experimental Setup
A.1 Data Processing
A.1.1 Pretraining data
@@ A.1.2 Alignment data
A.2 Model
A.2.1 Model architecture
@@ A.2.2 SE(3) diffusion
A.3 Training Details
A.3.1 Loss
@@ A.6 Tasks
Appendix B Additional Results

### v2

@@ 2.1 Divide and assembly diffusion framework
This spliced distance map $\bar{\mathbf{M}}$ initializes the edge representation, providing condition for the diffusion model, as shown in Appendix B.2.1. To enhance the spatial flexibility of domain assembly, we insert a 15-residue linker between adjacent domains and set its interaction with other resions to $\mathbf{-1}$. The model is tasked with learning the deformations generated by each domain in $\{\bar{\mathbf{S}}_{{D}_{1}},\bar{\mathbf{S}}_{{D}_{2}},...,\bar{\mathbf{S}}_{{D}_{m}}\}$ during the combination process as well as the patterns of the ultimate interactions.
To assemble generated domains, we utilize FrameDiff  (Yim et al., 2023), the $\mathrm{SE(3)}$ diffusion model to pretrain the domain assembly model. As illustrated in the Appendix B.2.1, the assembly module contains a series of folding blocks, where each folding block processes node representation, edge representation and structural frames.
@@ 2.2 Training
In this section, we provide the details of ProteinWeaver training, including datasets, pretraining, and preference alignments. More details can be found in Appendix B.2.1.
@@
Here, $\mathcal{L}_{\text{trans}}$ computes the loss between predicted translations with reference translations, while $\mathcal{L}_{\text{rot}}$ calculates the loss of rotation scores. $\mathcal{L}_{\text{atom}}$ represents the atom coordinate loss and $\mathcal{L}_{\text{pairwise}}$ computes the pairwise distance loss between predicted and reference positions. Following FrameDiff (Yim et al., 2023), we apply auxiliary losses when $t\leq{0.25}$. More details can be found in Appendix B.2.2.
@@ 3.1 Experimental setups
We evaluate ProteinWeaver’s performance using two tasks: (1) Domain assembly: this is a new protein backbone design task, introduced in our study, lacking existing deep learning-based baselines. It serves both as an appropriate scenario to evaluate our approach and as an important protein engineering task previously studied using traditional method (Huddy et al., 2024). (2) Backbone design: This task focuses on the design of de novo proteins, a subject of significant interest in the global research community. When generating backbone, we use the best of 3 operation (bo3). More details are provided in Appendix B.6.
@@
We pretrained the diffusion model for domain assembly using 5,835 filtered multi-domain PDBs from the RCSB database (Berman et al., 2000). For the preference alignment, we performed pairwise combinations of 100 single-domain structures, generating 10,000 pairs to build winner and loser datasets for alignment. Detailed methodology is provided in Appendix B.1.1 and B.1.2.
@@
We evaluate generation conformations as to their overall backbone quality, interface quality, novelty, and diversity. We also provide interface scTM and scRMSD metrics for domain assembly quality evaluation. More details are provided in Appendix B.5.
@@ 4.1 Diffusion modeling on protein backbone
Inspired by the considerable success of diffusion models across various fields (Ho et al., 2020; Song et al., 2020), researchers have applied this approach to protein structure and sequence design. Anand et al. pioneered this effort with a co-diffusion model for backbone, sequence, and sidechain generation using AlphaFold2’s Structure Module  (Anand & Achim, 2022) . Subsequent methods explored diffusion on inter-residue geometry  (Lee et al., 2023) and backbone dihedral angles  (Wu et al., 2024a). Current protein structure diffusion models primarily focus on end-to-end structure generation in SE3 or R3 space  (Yim et al., 2023), with extensions to function motif scaffolding  (Trippe et al., 2022; Yim et al., 2024). Chroma achieved more general conditioning and improved efficiency through geometrically constrained harmonic constraints  (Ingraham et al., 2023), while RFdiffusion demonstrated state-of-the-art designability with experimental validation  (Watson et al., 2023). Proteus (Wang et al., 2024) employs AlphaFold2’s graph-based triangle approach and multi-track interaction networks. It outperforms RFdiffusion in long-chain monomer generation and exceeds Chroma in complex structure generation. Also, sequence-structure co-design methods, such as MultiFlow (Campbell et al., 2024) and CarbonNovo (Ren et al., ) demonstrates superior designability compared to RFdiffusion.
@@ Conclusions.
While our framework demonstrates a new paradigm for protein backbone design, we acknowledge several technical aspects that warrant further investigation: (1) Recent advancements have revealed that flow matching-based protein backbone generation methods outperform diffusion-based approaches. An exciting avenue for future research lies in exploring whether replacing our current diffusion-based framework with a flow matching-based approach could yield further benefits; (2) Domain structure representation: Our current method explicitly represents local domain structures as distance maps. However, the efficacy of alternative pairwise representation methods remains an open question. (3) Integration of design stages: Currently, our approach employs two separate steps for backbone design. A promising direction for future research is developing an end-to-end unified framework that efficiently integrates these steps. Future research could further refine and extend this framework by addressing these questions, potentially revolutionizing our ability to create functional protein backbones with unprecedented precision.
Appendix A Brief Introduction of Domain Assembly
Domain assembly is a well-established task in protein design that has been extensively researched using traditional computational biology methods, such as RFdiffusion. To clarify, we provide the following examples that highlight the significance of this task:
[Domain-Assembly for Structural Design] A recent study demonstrated the use of helical protein blocks for designing extendable nanomaterials (Huddy et al., 2024). This work serves as a proof of concept for the domain-based design approach, but it focused solely on pure helical-bundle assemblies. In contrast, our study generalizes this idea to encompass a broader range of domains with structural variations. We present the first application of deep learning methods for general domain assembly, marking an important new paradigm in this field.
Domain assembly for functional design] Nearly two decades ago, research showed that well-assembled domains, through interface-directed evolution, could dramatically enhance both affinity and specificity—achieving over 500-fold and 2,000-fold increases, respectively (Huang et al., 2008). More recently, another study revealed that assembling a substrate recruitment domain to an enzyme significantly improved its functional specificity (Park et al., 2023). These studies underscore the validity and importance of our motivation for pursuing domain-assembly-based protein structure and function design, a concept we introduced in our manuscript. We have added more references in the revised manuscript to emphasize this importance.
Appendix B Experimental Setup
B.1 Data Processing
B.1.1 Pretraining data
@@ B.1.2 Alignment data
B.2 Model
B.2.1 Model architecture
@@ B.2.2 SE(3) diffusion
B.3 Training Details
B.3.1 Loss
@@ B.6 Tasks
Appendix C Additional Results
@@
Figure 11: We clustered natural domains in the CATH dataset based on their topological differences, resulting in approximately 530 domain structures that represent a wide distribution of protein domains. This dataset allows us to evaluate the effects of structural variations effectively. We performed pairwise assemblies of these domains and assessed the quality of the designed structures using the scTM score. Our analysis included a comparison of secondary structure ratios at the assembly interface, which we believe directly affects domain assembly.
@@
Table 4: Comparison between ProteinWeaver and co-design models evaluated using various lengths ranging from 100 to 800. For each length, we randomly sampled 50 native PDB structures from RCSB as a golden reference for the task. The reported results are the mean±std of repetitive experiments. When the length is 100, the current co-design model generates too few multi-domain proteins and is not statistically significant, so it is not reported.
length100 length200
Multiflow 0.96±0.04 1.10±0.71 0.71±0.08 0.33 0.95±0.04 1.61±1.73 0.90±0.03 * 0.71±0.07 0.42
CarbonNovo 0.91±0.14 1.16±1.03 0.69±0.09 0.71 0.94±0.09 1.18±1.47 0.97±0.01 * 0.71±0.08 0.50
length300 length500
Multiflow 0.96±0.06 2.14±3.24 0.91±0.04 * 0.71±0.06 0.58 0.83±0.10 8.48±5.32 0.84±0.07 * 0.68±0.06 0.67
CarbonNovo 0.95±0.08 1.33±1.59 0.93±0.11 * 0.74±0.05 0.31 0.85±0.15 4.07±4.14 0.83±0.17 * 0.68±0.05 0.67
length600 length800
Multiflow 0.61±0.13 12.41±4.74 0.49±0.11 * 0.71±0.07 0.62 0.37±0.07 25.86±3.18 0.27±0.06 * 0.54
CarbonNovo 0.87±0.09 4.20±4.09 0.81±0.09 * 0.70±0.06 0.93 0.52±0.13 16.53±5.39 0.57±0.17 * 0.67±0.03 1.00
Table 5: The scTM calculation involves generating eight sequences for each designed backbone using proteinMPNN, followed by structure prediction with ESMFold for each sequence. The sequence with the highest scTM score is then selected for further analysis. As both the number of sampled sequences and the length of the designed proteins increase, the computational demands rise significantly, as illustrated in the table below.
length seq_per_sample (unit: second)
1 2 4 8
50 9.9 12.3 16.3 23
100 10.9 12 17.2 25
200 16 21.3 32 52
300 30 41 66 113
400 51 76 132 216
500 89 130 211 377
Table 6: Evaluation of Preference Alignment Methods. We evaluated the performance of SPPO with different preference alignment methods: SFT (Supervised Fine-Tuning) and SFT + DPO (Direct Preference Optimization). To ensure the robustness of our implementation, we tested DPO with various beta parameters. The beta parameter controls the degree of difference between the fine-tuned model and the reference model. A larger beta value makes it less likely for the model to deviate from the reference model during the fine-tuning process.
length100 length200
scTM ↑ scRMSD ↓ Max TM ↓ Max Clust ↑ scTM ↑ scRMSD ↓ Max TM ↓ Max Clust ↑
SPPO 0.91±0.14 1.27±1.84 0.61±0.07 0.59 0.88±0.13 2.47±3.63 0.65±0.07 0.67
SFT 0.91±0.12 1.20±1.79 0.62±0.06 0.61 0.86±0.12 2.61±2.44 0.67±0.06 0.68
SFT+DPO (beta=10) 0.90±0.11 1.45±1.96 0.62±0.04 0.62 0.87±0.13 2.57±2.41 0.67±0.06 0.68
SFT+DPO (beta=1.0) 0.88±0.14 1.89±2.18 0.62±0.05 0.62 0.83±0.13 2.93±2.41 0.68±0.07 0.68
SFT+DPO (beta=0.1) 0.88±0.14 1.89±2.18 0.63±0.05 0.62 0.78±0.14 3.74±2.83 0.67±0.05 0.72
length300 length500
scTM ↑ scRMSD ↓ Max TM ↓ Max Clust ↑ scTM ↑ scRMSD ↓ Max TM ↓ Max Clust ↑
SPPO 0.88±0.13 2.47±3.63 0.67±0.06 0.86 0.82±0.10 4.39±2.72 0.67±0.07 0.78
SFT 0.86±0.10 3.15±2.20 0.68±0.06 0.85 0.72±0.14 8.30±3.19 0.72±0.07 0.89
SFT+DPO (beta=10) 0.85±0.09 3.24±2.37 0.68±0.06 0.85 0.72±0.13 8.26±3.05 0.72±0.07 0.89
SFT+DPO (beta=1.0) 0.86±0.10 3.02±2.05 0.68±0.06 0.85 0.77±0.13 5.79±3.66 0.72±0.06 0.61
SFT+DPO (beta=0.1) 0.80±0.12 4.10±2.91 0.70±0.03 0.87 0.48±0.09 12.99±3.04 0.78±0.02 0.97
length600 length800
scTM ↑ scRMSD ↓ Max TM ↓ Max Clust ↑ scTM ↑ scRMSD ↓ Max TM ↓ Max Clust ↑
SPPO 0.72±0.15 7.15±2.74 0.68±0.07 0.75 0.63±0.13 9.17±5.81 0.68±0.06 0.73
SFT 0.66±0.16 9.71±5.61 0.69±0.05 0.79 0.28±0.16 31.73±13.14 0.67±0.04 0.95
SFT+DPO (beta=10) 0.68±0.14 9.40±4.91 0.69±0.05 0.79 0.34±0.22 27.11±11.90 0.68±0.036 0.93
SFT+DPO (beta=1.0) 0.67±0.15 9.24±4.72 0.71±0.06 0.68 0.55±0.10 13.89±4.54 0.70±0.03 0.85
SFT+DPO (beta=0.1) 0.40±0.07 20.24±12.88 0.76±0.03 1.00 0.27±0.10 33.50±15.72 1.00
Table 7: We explored the impact of varying reward assignments by adjusting preferences from scTM to interface scTM, as well as combining both metrics. For the combined approach, we selected the sample that exhibited the best performance in both scTM and interface scTM. We believe that scTM serves as a general quality metric for backbone structures, while interface scTM focuses on optimizing inter-domain interactions.
length100 length200
scTM ↑ scRMSD ↓ Intf. scTM ↑ Max TM ↓ Max Clust ↑ scTM ↑ scRMSD ↓ Intf. scTM ↑ Max TM ↓ Max Clust ↑
ProteinWeaver (scTM) w/o bo3 0.91±0.14 1.27±1.84 0.75±0.14 0.61±0.07 0.59 0.88±0.13 2.47±3.63 0.73±0.14 0.65±0.07 0.67
ProteinWeaver (interface scTM) w/o bo3 0.86±0.11 1.73±1.74 0.60±0.12 0.62±0.08 0.61 0.84±0.12 2.69±2.56 0.71±0.15 0.66±0.05 0.69
ProteinWeaver (scTM + interface scTM) w/o bo3 0.91±0.08 1.12±0.81 0.77±0.14 0.61±0.06 0.60 0.88±0.12 2.14±2.07 0.76±0.14 0.65±0.06 0.67
length300 length500
scTM ↑ scRMSD ↓ Intf. scTM ↑ Max TM ↓ Max Clust ↑ scTM ↑ scRMSD ↓ Intf. scTM ↑ Max TM ↓ Max Clust ↑
ProteinWeaver (scTM) w/o bo3 0.88±0.13 2.47±3.63 0.74±0.07 0.67±0.06 0.86 0.82±0.10 4.39±2.72 0.70±0.14 0.67±0.07 0.78
ProteinWeaver (interface scTM) w/o bo3 0.82±0.12 3.79±2.89 0.66±0.17 0.66±0.06 0.83 0.79±0.14 5.23±3.89 0.67±0.15 0.69±0.08 0.79
ProteinWeaver (scTM + interface scTM) w/o bo3 0.86±0.10 2.66±2.21 0.74±0.14 0.66±0.06 0.88 0.84±0.12 4.21±3.33 0.72±0.14 0.68±0.06 0.78
length600 length800
scTM ↑ scRMSD ↓ Intf. scTM ↑ Max TM ↓ Max Clust ↑ scTM ↑ scRMSD ↓ Intf. scTM ↑ Max TM ↓ Max Clust ↑
ProteinWeaver (scTM) w/o bo3 0.72±0.15 7.15±2.74 0.70±0.14 0.68±0.07 0.75 0.63±0.13 9.17±5.81 0.59±0.14 0.68±0.06 0.73
ProteinWeaver (interface scTM) w/o bo3 0.69±0.14 8.79±5.14 0.61±0.17 0.67±0.07 0.78 0.54±0.11 14.36±4.35 0.52±0.13 0.70±0.09 0.88
ProteinWeaver (scTM + interface scTM) w/o bo3 0.72±0.13 7.28±4.72 0.68±0.15 0.67±0.06 0.76 0.62±0.09 9.34±4.88 0.57±0.15 0.68±0.08 0.75
Table 8: We have also experimented with adding triangular attention to ProteinWeaver and evaluated its performance. We replaced the edge transition from MLP to triangular attention, and the performance results are shown in below.
Model length100 length200
scTM scRMSD Interface scTM diversity novelty scTM scRMSD Interface scTM diversity novelty
ProteinWeaver(MLP) 0.90±0.14 1.26±1.97 0.73±0.07 0.60 0.61±0.07 0.86±0.11 2.59±3.70 0.71±0.15 0.68 0.66±0.06
ProteinWeaver(Triangular attention) 0.92±0.11 1.66±1.70 0.76±0.12 0.58 0.62±0.10 0.88±0.10 1.99±1.47 0.77±0.12 0.67 0.67±0.06
length300 length500
scTM scRMSD Interface scTM diversity novelty scTM scRMSD Interface scTM diversity novelty
ProteinWeaver(MLP) 0.86±0.10 3.16±2.40 0.71±0.16 0.86 0.67±0.06 0.78±0.14 5.77±4.36 0.68±0.15 0.76 0.67±0.06
ProteinWeaver(Triangular attention) 0.88±0.09 2.64±1.96 0.73±0.13 0.85 0.70±0.07 0.80±0.12 5.35±3.96 0.68±0.19 0.72 0.68±0.06
length600 length800
scTM scRMSD Interface scTM diversity novelty scTM scRMSD Interface scTM diversity novelty
ProteinWeaver(MLP) 0.69±0.18 8.79±5.74 0.69±0.15 0.76 0.69±0.06 0.54±0.12 12.87±4.87 0.58±0.09 0.73 0.67±0.07
ProteinWeaver(Triangular attention) 0.69±0.19 8.52±5.56 0.66±0.12 0.76 0.68±0.08 0.56±0.11 12.18±4.43 0.57±0.09 0.73 0.67±0.07